# The drastic impact of Eddington-limit induced mass ejections on massive star populations

### v1

@@ 4.1 Neglecting Eddington-limit induced mass ejections
Figure 9: HRD of the LMC population (left), luminosity distributions of the WR subtypes (center), and luminosity distribution of the cool supergiants (right) compared to the model predictions of our Model Binary.
@@ 4.2.1 The OB and VMS population
An HRD comparing the synthetic binary population at LMC metallicity with observations is shown in the left panel of Fig. 9. Overall, it is quite similar to the single-star population (Fig.3), apart from the appearance of a He-star population located below the WR stars. The predicted number of O stars is about 3200, and thus a factor of 1.2 times higher than in the single-star models. This number is sensitive to both the assumed post-interaction evolution of primaries, secondaries, and merger products, as well as the way one counts the stars that are seen by an observer. Given that the difference remains within a few tens of percent, we consider our models to still provide a good representation of the LMC population. Interestingly, because the model weights are assigned according to the IMF of the primary mass and the secondaries are assumed to follow a uniform mass-ratio distribution, the inclusion of binaries reduces the predicted number of VMS to 19, bringing the models at the high mass end into closer agreement with observations.
@@ 4.2.2 The WR population
For WC/WO stars, our model predicts 16 objects compared to the 26 observed. Since we are only calculating our models until core He depletion, it might be that we accidentally cut off a part of this population. However, given that core C burning takes only about $10\%$ of the time a star spends core He burning, this would marginally increase the numbers and might only help to explain the faint WO stars seen in the LMC.
@@ 4.2.3 The cool supergiant population
Binary interaction and binary evolution are often associated with mass transfer and is considered as a shortcut to form WR stars, at the expense of skipping the RSG phase and thus leading to too few RSG in a binary population. In contrast to these expectations, our binary population contains 270 RSGs, which is slightly more stars than in the single-star population. The RSGs in our population originate both from mergers and ejected secondaries. This can be seen in the right panel of Fig. 9, which illustrates the luminosity distribution of the cool supergiant population, separated into contributions from non-interacting binaries, secondaries, main-sequence mergers, and post-main-sequence mergers. The fraction of RSGs originating from mergers decreases with increasing luminosity: while $\approx 25\%$ of the RSGs at $\log(L/L_{\odot})\lesssim 5.4$ are merger products, this fraction drops sharply below $5\%$ at higher luminosities. This behavior reflects the stability criteria for mass transfer adopted in our binary grids: lower-mass donors often undergo unstable mass transfer, whereas mass transfer from higher-mass donors is typically stable, and thus leading to fewer mergers at higher initial masses.
@@ 5 Summary
In this work, we present a new mass-loss prescription to account for mass ejections during the evolution of luminous massive stars approaching the Eddington limit, such as those expected during an LBV phase. Within our models, the mass ejections are triggered when stellar models inflate their envelope as a consequence of reaching the Eddington limit beyond a threshold value. The HD limit and the WR populations of the LMC and SMC are used as benchmarks to calibrate the maximum inflation a star can have before going through a mass ejection episode, effectively removing inflation.
@@
From our models, Eddington-limit induced mass ejections can be triggered during two main phases. Stars more massive than $M_{\mathrm{ini}}\gtrsim 50\,M_{\odot}$ enter an LBV-like phase during the main sequence. Before the mass ejection episode, they have mass-loss rates comparable to those of the early-O supergiants, while after the mass ejections, they have mass-loss rates comparable to the WNh stars, both observed in the same region of the HRD. Stars with $M_{\mathrm{ini}}\approx 25\,M_{\odot}\text{\,--\,}40\,M_{\odot}$ evolve first into RSGs. As soon as they have expelled their H-rich envelope, radiation pressure will dominate in the envelope, associated with envelope inflation. After the RSG phase, these stars experience short mass ejections in a region similar to those of the faintest observed LBVs.
@@
Our new Eddington-limit induced mass ejection model exhibits only a slight $Z$ dependence, as the opacity governs how much a star inflates. This $Z$ dependence is evident when comparing the LMC and SMC populations, although it remains consistent with observed data. Due to the weak $Z$ dependence, one would expect that still at very low metallicities, only a few stars reside beyond the HD limit, observable as cool supergiants rather than RSGs.
@@
Despite its simplicity, the models using our new physical setup successfully replicate many of the features observed in stellar populations, even at low $Z$, and provide a definitive improvement compared to models without Eddington-limit induced mass ejections. Within the models, the Eddington-limit induced mass ejections occur in regions of the HRD where LBVs are observed. These mass ejections explain the lack of stars beyond the HD limit, including the upper limit of RSG luminosities. Furthermore, the number of observed O-type stars, WRs, and RSGs in the LMC and SMC can be matched. When using strong $Z$-independent RSG winds such as those from yan1:23, we can explain the faintest observed single WR stars, but only when using the Eddington-limit induced mass ejections, we can reproduce the observed low-$Z$ WO-type stars in the SMC and IC 1613.
@@
We combined our single-star models with large, detailed grids of binary evolution models at LMC metallicity. According to our binary population, binary interactions still play a fundamental role in forming WR stars, especially for the H-free WN stars. However, a large fraction of the WR stars can be formed by stars in non-interacting binaries and mergers. One of the limitations of our Eddington-limit induced mass ejection model is that all stars above $\gtrsim 150\,M_{\odot}$ converge to the same track, leading to an excess of luminous H-free WR stars. Our binary population assumes that all stars are born in binary systems, with 70% of the O-stars found in close binaries where they can interact. Interestingly, our new model predicts that less than $\sim 40\%$ of the WR stars are in close binary systems, being in agreement with observations. This further underlines the importance of Eddington-limit induced mass ejections in stellar populations.
@@ Appendix A Additional information on the assumed input physics
To properly model the nuclear reactions of massive stars, we use a nuclear network inspired by the stellar evolution code STERN (heg1:00; jin1:24). We include 53 reaction rates for the p-p chains, CNO, Ne-Na, and Mg-Al burning, which are important already during core hydrogen and helium burning in VMS.
@@ B.2 Binary star population
Similar to before, we can use the total mass enclosed in the stars to calculate the normalization factor by
@@ Appendix C Detailed definitions of the distinct stellar groups
WR stars exhibit optically thick winds, which lead to strong emission lines in their spectra. These emission features in the observations are used for classification. However, stellar evolution models do not output the spectrum of a star. To identify when a stellar model enters the WR phase, we follow the method described agu1:20 and calculate the optical depth of a stellar model using
@@
WR stars have multiple subtypes, each characterized by emission lines from different elements, namely nitrogen, carbon, and oxygen. These subtypes are associated with distinct evolutionary phases and serve as excellent benchmarks for comparison of model predictions with observed populations. In this work, we define the following categories for the nitrogen-rich WN-type stars: H-rich WNs that have $X_{\mathrm{H}}>0.4$, H-poor WNs with $0.05<X_{\mathrm{H}}\leq 0.4$, and H-free WNs are those WR stars with $X_{\mathrm{H}}\leq 0.05$ and $X_{\mathrm{C}}<0.005$. While observational WC and WO-type stars are classified by their different strengths in carbon and oxygen lines, recent studies indicate that the different emission strengths are due to a temperature effect rather than an abundance effect (aad1:22; san2:25). Therefore, we do not differentiate between these two classes, and within our models, we define the WC/WO phase as soon as $X_{\mathrm{C}}\geq 0.005$.
@@
As VMS, we define all stars with a luminosity above $\log(L/L_{\odot})\geq 6.5$ (i.e., $M_{\mathrm{ini}}\gtrsim 150\,M_{\odot}$). These stars, even though extremely rare in stellar populations, are dominating the integrated spectra of actively star-forming galaxies and are therefore of particular interest when trying to mimic real stellar populations. Note that the VMS are also included in the WR and O-type stars sample.
@@ D.1 Changing the RSG mass-loss rates
Figure 10: HRDs of the massive star population (symbols) in the LMC (left), and SMC (right), compared to the synthetic population of the Model RSG (contours). The symbols and colors have the same meaning as in Fig. 1.
One of the biggest uncertainties in modeling massive star evolution, besides the Eddington-limit induced mass ejections, is the mass lost via the winds of RSG. Comparing RSG mass-loss rates derived using different model assumptions (such as the launching mechanism of the wind) shows that the measured RSG mass-loss rate can differ by up to two orders of magnitude (e.g., dec1:21; yan1:23; bea1:23; ant1:24; dec1:24). These differences arise partly from the method of modeling itself (e.g. see figure 13 in ant1:24). When comparing different recent mass-loss prescriptions (e.g. bea1:23; yan1:23; dec1:24), it becomes evident that most of them align within a factor of a few for luminosities above $\log(L/L_{\odot})\gtrsim 5$ (i.e., the potential progenitors of WR stars) and that stronger mass-loss rates compared to the “classic” nie1:90 are observed. Still in many publicly available stellar evolution grids, the mass-loss recipe of nie1:90 is commonly employed, sometimes even with a $Z$ dependence. Here, we present our Model RSG set, where we have switched from the yan1:23 mass-loss recipe to that of nie1:90, but without any $Z$ dependence.
@@ D.2 More efficient semiconvective mixing
The luminosity distribution of cool supergiants for this model is shown in Fig. 8. While the blue-loop evolution successfully populates both the BSG and YSG regions, most of the predicted stars are found at $\log(L/L_{\odot})\gtrsim 5.5$, a regime where observations show few to none objects. Moreover, the Model sc10 predicts more YSGs than BSGs, opposite to what is observed. Both discrepancies may stem from our adopted semiconvection and overshooting parameters. As shown by sch1:19, stars at SMC metallicity only undergo blue loops for specific combinations of overshooting and semiconvection. It is still unclear whether these mixing parameters, such as overshooting, may vary with mass and what this implies for the cool supergiant population (e.g., cla1:16; jer1:18; gil1:19).
@@
Figure 11: HRDs of the massive star population (symbols) in the LMC (left), and SMC (right), compared to the synthetic population of the Model sc10 (contours). The symbols and colors have the same meaning as in Fig. 1.
@@ Appendix E Additional Figures
Figure 12: HRDs showing the tracks of our Model Eject grid for LMC (left) and SMC metallicity (right). The Eddington-limit induced mass ejection episodes of our models are indicated by bold gray outlines around the tracks. Black dots on the tracks indicate equidistant timesteps of $\Delta t=$30\,000\text{\,}\mathrm{y}\mathrm{r}$\$. For comparison, the known LBVs of the LMC and SMC are shown as triangles during quiescence (left triangle) and outburst (right triangle) (hum1:16; kal1:18).

### v2

@@ 4.1 Neglecting Eddington-limit induced mass ejections
Figure 9: Hertzsprung-Russell diagram of the LMC population (left), luminosity distributions of the WR subtypes (center), and luminosity distribution of the cool supergiants (right) compared to the model predictions of our Model Binary.
@@ 4.2.1 The OB and VMS population
An HRD comparing the synthetic binary population at LMC metallicity with observations is shown in the left panel of Fig. 9. Overall, it is quite similar to the single-star population (Fig.3), apart from the appearance of a He-star population located below the WR stars. The predicted number of O stars is about 3200, a factor of 1.2 times higher than in the single-star models. This number is sensitive to both the assumed post-interaction evolution of primaries, secondaries, and merger products, as well as the way one counts the stars that are seen by an observer. Given that the difference remains within a few tens of percent, we consider our models to still provide a good representation of the LMC population. Interestingly, because the model weights are assigned according to the IMF of the primary mass and the secondaries are assumed to follow a uniform mass-ratio distribution, the inclusion of binaries reduces the predicted number of VMSs to 19, bringing the models at the high mass end into closer agreement with observations.
@@ 4.2.2 The WR population
For WC/WO stars, our model predicts 16 objects compared to the 26 observed. Since we are only calculating our models until core He depletion, it might be that we accidentally cut off a part of this population. However, given that core C burning takes only about $10\%$ of the time a star spends core He burning, this would marginally increase the numbers and might only help explain the faint WO stars seen in the LMC.
@@ 4.2.3 The cool supergiant population
Binary interaction and binary evolution are often associated with mass transfer, which is considered a shortcut to forming WR stars at the expense of skipping the RSG phase and thus causes there to be fewer RSGs in a binary population. In contrast to these expectations, our binary population contains 270 RSGs, which is slightly more stars than in the single-star population. The RSGs in our population originate both from mergers and ejected secondaries. This can be seen in the right panel of Fig. 9, which illustrates the luminosity distribution of the cool supergiant population, separated into contributions from non-interacting binaries, secondaries, main-sequence mergers, and post-main-sequence mergers. The fraction of RSGs originating from mergers decreases with increasing luminosity: while $\approx 25\%$ of the RSGs at $\log(L/L_{\odot})\lesssim 5.4$ are merger products, this fraction drops sharply below $5\%$ at higher luminosities. This behavior reflects the stability criteria for mass transfer adopted in our binary grids: lower-mass donors often undergo unstable mass transfer, whereas mass transfer from higher-mass donors is typically stable, thus leading to fewer mergers at higher initial masses.
@@ 5 Summary
In this work, we have presented a new mass-loss prescription to account for mass ejections during the evolution of luminous massive stars approaching the Eddington limit, such as those expected during an LBV phase. Within our models, the mass ejections are triggered when stellar models inflate their envelope as a consequence of reaching the Eddington limit beyond a threshold value. The HD limit and the WR populations of the LMC and SMC were used as benchmarks to calibrate the maximum inflation a star can have before going through a mass ejection episode, thus effectively removing inflation.
@@
From our models, Eddington-limit induced mass ejections can be triggered during two main phases. Stars more massive than $M_{\mathrm{ini}}\gtrsim 50\,M_{\odot}$ enter an LBV-like phase during the main sequence. Before the mass ejection episode, they have mass-loss rates comparable to those of the early O supergiants, while after the mass ejections, they have mass-loss rates comparable to the WNh stars; both were observed in the same region of the HRD. Stars with $M_{\mathrm{ini}}\approx 25\,M_{\odot}\text{\,--\,}40\,M_{\odot}$ evolve first into RSGs. As soon as they expel their H-rich envelope, radiation pressure dominates in the envelope, which is associated with envelope inflation. After the RSG phase, these stars experience short mass ejections in a region similar to those of the faintest observed LBVs.
@@
Our new Eddington-limit induced mass ejection model exhibits only a slight $Z$ dependence, as the opacity governs how much a star inflates. This $Z$ dependence is evident when comparing the LMC and SMC populations, although it remains consistent with observed data. Due to the weak $Z$ dependence, one would still expect that at very low metallicities, only a few stars would reside beyond the HD limit. These stars would be observable as cool supergiants rather than RSGs.
@@
Despite their simplicity, the models using our new physical setup successfully replicate many of the features observed in stellar populations, even at low $Z$, and they provide a definitive improvement compared to models without Eddington-limit induced mass ejections. Within the models, the Eddington-limit induced mass ejections occur in regions of the HRD where LBVs are observed. These mass ejections explain the lack of stars beyond the HD limit, including the upper limit of RSG luminosities. Furthermore, the number of observed O-type stars, WRs, and RSGs in the LMC and SMC can be matched. When using strong $Z$-independent RSG winds, such as those from yan1:23, we can explain the faintest observed single WR stars; however, it is only when we use the Eddington-limit induced mass ejections that we can reproduce the observed low-$Z$ WO-type stars in the SMC and IC 1613.
@@
We combined our single-star models with large detailed grids of binary evolution models at LMC metallicity. According to our binary population, binary interactions still play a fundamental role in forming WR stars, especially for the H-free WN stars. However, a large fraction of the WR stars can be formed by stars in non-interacting binaries and mergers. One of the limitations of our Eddington-limit induced mass ejection model is that all stars $\gtrsim 150\,M_{\odot}$ converge to the same track, leading to an excess of luminous H-free WR stars. Our binary population assumes that all stars are born in binary systems, with 70% of the O-stars found in close binaries where they can interact. Interestingly, our new model predicts that about less than $40\%$ of the WR stars are in close binary systems, which is in agreement with observations. This further underlines the importance of Eddington-limit induced mass ejections in stellar populations.
@@ Appendix A Additional information on the assumed input physics
To properly model the nuclear reactions of massive stars, we use a nuclear network inspired by the stellar evolution code STERN (heg1:00; jin1:24). We include 53 reaction rates for the p-p chains, CNO, Ne-Na, and Mg-Al burning, which are important already during core hydrogen and helium burning in VMSs.
@@ B.2 Binary star population
Similar to the calculation of the single-star population, we can use the total mass enclosed in the stars to calculate the normalization factor by
@@ Appendix C Detailed definitions of the distinct stellar groups
Wolf-Rayet stars exhibit optically thick winds, which lead to strong emission lines in their spectra. These emission features in the observations are used for classification. However, stellar evolution models do not output the spectrum of a star. To identify when a stellar model enters the WR phase, we follow the method described agu1:20 and calculate the optical depth of a stellar model using
@@
Wolf-Rayet stars have multiple subtypes, each characterized by emission lines from different elements, namely nitrogen, carbon, and oxygen. These subtypes are associated with distinct evolutionary phases and serve as excellent benchmarks for comparison of model predictions with observed populations. In this work, we define the following categories for the nitrogen-rich WN-type stars: H-rich WNs that have $X_{\mathrm{H}}>0.4$, H-poor WNs with $0.05<X_{\mathrm{H}}\leq 0.4$, and H-free WNs are those WR stars with $X_{\mathrm{H}}\leq 0.05$ and $X_{\mathrm{C}}<0.005$. While observational WC and WO-type stars are classified by their different strengths in carbon and oxygen lines, recent studies indicate that the different emission strengths are due to a temperature effect rather than an abundance effect (aad1:22; san2:25). Therefore, we do not differentiate between these two classes, and within our models, we define the WC/WO phase as soon as $X_{\mathrm{C}}\geq 0.005$.
@@
As VMSs, we define all stars with a luminosity above $\log(L/L_{\odot})\geq 6.5$ (i.e., $M_{\mathrm{ini}}\gtrsim 150\,M_{\odot}$). These stars, even though extremely rare in stellar populations, are dominating the integrated spectra of actively star-forming galaxies and are therefore of particular interest when trying to mimic real stellar populations. Note that the VMSs are also included in the WR and O-type stars sample.
@@ D.1 Changing the RSG mass-loss rates
Figure 10: Hertzsprung-Russell diagrams of the massive star population (symbols) in the LMC (left), and SMC (right), compared to the synthetic population of the Model RSG (contours). The symbols and colors have the same meaning as in Fig. 1.
One of the biggest uncertainties in modeling massive star evolution, besides the Eddington-limit induced mass ejections, is the mass lost via the winds of RSG. Comparing RSG mass-loss rates derived using different model assumptions (such as the launching mechanism of the wind) shows that the measured RSG mass-loss rate can differ by up to two orders of magnitude (e.g., dec1:21; yan1:23; bea1:23; ant1:24; dec1:24). These differences arise partly from the method of modeling itself (e.g., see figure 13 in ant1:24). When comparing different recent mass-loss prescriptions (e.g., bea1:23; yan1:23; dec1:24), it becomes evident that most of them align within a factor of a few for luminosities above $\log(L/L_{\odot})\gtrsim 5$ (i.e., the potential progenitors of WR stars) and that stronger mass-loss rates compared to the “classic” nie1:90 are observed. Still in many publicly available stellar evolution grids, the mass-loss recipe of nie1:90 is commonly employed, sometimes even with a $Z$ dependence. Here, we present our Model RSG set, where we have switched from the yan1:23 mass-loss recipe to that of nie1:90, but without any $Z$ dependence.
@@ D.2 More efficient semiconvective mixing
The luminosity distribution of cool supergiants for this model is shown in Fig. 8. While the blue-loop evolution successfully populates both the BSG and YSG regions, most of the predicted stars are found at $\log(L/L_{\odot})\gtrsim 5.5$, a regime where observations show few to no objects. Moreover, the Model sc10 predicts more YSGs than BSGs, opposite to what is observed. Both discrepancies may stem from our adopted semiconvection and overshooting parameters. As shown by sch1:19, stars at SMC metallicity only undergo blue loops for specific combinations of overshooting and semiconvection. It is still unclear whether these mixing parameters, such as overshooting, may vary with mass and what this implies for the cool supergiant population (e.g., cla1:16; jer1:18; gil1:19).
@@
Figure 11: Hertzsprung-Russell diagrams of the massive star population (symbols) in the LMC (left) and SMC (right) compared to the synthetic population of the Model sc10 (contours). The symbols and colors have the same meaning as in Fig. 1.
@@ Appendix E Additional figures
Figure 12: Hertzsprung-Russell diagrams showing the tracks of our Model Eject grid for LMC (left) and SMC metallicity (right). The Eddington-limit induced mass ejection episodes of our models are indicated by bold gray outlines around the tracks. Black dots on the tracks indicate equidistant timesteps of $\Delta t=$30\,000\text{\,}\mathrm{y}\mathrm{r}$\$. For comparison, the known LBVs of the LMC and SMC are shown as triangles during quiescence (left triangle) and outburst (right triangle) (hum1:16; kal1:18).